OpenStudy (anonymous):

Solve the equation one-half times x plus three-fourths equals one-fourth for x. x equals one-fourth x equals negative one-fourth x equals 1 x equals negative 1

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

Wait, is that of x or for x i dont understand that part i may be wrong therefore.

OpenStudy (anonymous):

I just solved it as x/2 + 3/4 = 1/4 you get 2x+3 = 4/4 2x=-2 x = -1

OpenStudy (anonymous):

1/2x + 3/4 = 1/4 --- subtract 3/4 from both sides 1/2x = 1/4 - 3/4 1/2x = - 1/2 -- divide by 1/2 x = (-1/2) / (1/2) x = - 1